how do solve -3/5[2/3-1/2(3/4-1/2)]
Answer by tommyt3rd(5050)

-3/5[2/3-1/2(3/4-1/2)]=
(-3/5)(2/3)-(-3/5)(1/2)(3/4-1/2)=
-2/5+(3/10)(1/4)=
-2/5+3/40=
-16/40+3/40=
-13/40